OpenStudy (jocelynstalnaker):

Number 2 and letter E.

OpenStudy (jocelynstalnaker):

2. While in the United States visiting her grandmother, Kendra copied the famous family apple pie recipe. When she returns to England, she plans on making the same apple pie for her friends. To do this, Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams. -A. Write a function that shows grams as a function of cups. Use g(x). -B. Given this scenario, determine a reasonable domain for this function. -C. Graph the relationship between the number of cups and grams of flour. -D. Evaluate g(x) for any value of x within the domain. Interpret the meaning of your solution in the context of this problem. -E. Look at the method you used to evaluate your function in part D. Explain another way you can evaluate the function for the same input. -F. Interpret what it means to have g(-2), and determine if this is reasonable given the situation.

OpenStudy (jocelynstalnaker):

@goformit100

OpenStudy (goformit100):

are both questions attached ?

OpenStudy (jocelynstalnaker):

This is just one question, I just need help with D-

OpenStudy (goformit100):

Do you know the equation, or I create a equation for it ?

OpenStudy (goformit100):

Let me read it again : Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams. Evaluate g(x) for any value of x within the domain. Interpret the meaning of your solution in the context of this problem.

OpenStudy (goformit100):

f(x) = 0.5x + 2x = 64 + 256 i think this will be the equation.

OpenStudy (jocelynstalnaker):

Wait wouldnt that one be for A?

OpenStudy (goformit100):

ya....

OpenStudy (goformit100):

Question D as you suggested is the continued question of A. in this question D you have to put a specific value in place of x. Nothing more. Simply give x a value and solve the equation formed.

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

Should I be more clear ?

OpenStudy (jocelynstalnaker):

Yes, please.

OpenStudy (goformit100):

did you get the same equation for question A ?

OpenStudy (jocelynstalnaker):

Yes.

OpenStudy (goformit100):

Was it correct ?

OpenStudy (jocelynstalnaker):

Ok, wait, I'm confused. Let me just start over with this whole problem.

OpenStudy (goformit100):

Sure

OpenStudy (jocelynstalnaker):

So 0.5 cups is 64 grams and 2 cups is 256 grams.

OpenStudy (jocelynstalnaker):

So f(x)= 0.5x+64 and f(x)=2x+256

OpenStudy (goformit100):

Let me read it again : Kendra needs to convert the flour from cups to grams. She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams. Evaluate g(x) for any value of x within the domain. Interpret the meaning of your solution in the context of this problem.

OpenStudy (goformit100):

i think equation should be of the form : - like, 1 km = 1000 m do you agree ? then, x km = 1000x m

OpenStudy (jocelynstalnaker):

Yes. LIke, 0.5cups(has a mass of) = 64grams

OpenStudy (jocelynstalnaker):

Like*

OpenStudy (goformit100):

0.5 cups = 64 gram 0.5x = 64x will this work

OpenStudy (jocelynstalnaker):

No because x cant be cups and grams.

OpenStudy (goformit100):

@ParthKohli Urgent help

OpenStudy (jocelynstalnaker):

Im so confused..

OpenStudy (goformit100):

Don't worry @JocelynStalnaker Now @ParthKohli is there to help us to confused person.... jk

OpenStudy (goformit100):

*two

OpenStudy (goformit100):

Put 5 in place of @ParthKohli 's told equation.

ganeshie8 (ganeshie8):

@JocelynStalnaker are you done with parts A,B,C,D ?

OpenStudy (jocelynstalnaker):

Yes

OpenStudy (jocelynstalnaker):

And i have F

OpenStudy (jocelynstalnaker):

Just not E.

ganeshie8 (ganeshie8):

May I see your work quick... because E depends on what you did in the previous parts..

OpenStudy (jocelynstalnaker):

A- g(x)=128x B- 5 can be a reasonable domain. C- A graph-- D- g(5)=128(5). g(5)=640.

ganeshie8 (ganeshie8):

Excellent ! everything looks good except for part B

OpenStudy (jocelynstalnaker):

Whats wrong with B?

ganeshie8 (ganeshie8):

lets work part E & F first, may be after that we will get back to part B. okay ?

OpenStudy (jocelynstalnaker):

Ok

ganeshie8 (ganeshie8):

so we need to find another way to find the value of g(x) when x = 5

OpenStudy (jocelynstalnaker):

Yes, But how would i do that?

ganeshie8 (ganeshie8):

We're given below points : `She knows that 0.5 cup of flour has a mass of 64 grams and 2 cups of flour has a mass of 256 grams.` pick any one point and set up a proportion

ganeshie8 (ganeshie8):

2 cups = 256 grams 5 cups = ? grams

OpenStudy (jocelynstalnaker):

640

ganeshie8 (ganeshie8):

You need to show the work also, otherwise it will be hard for your teacher to see any difference between methods in part D and E \(\large \dfrac{2}{5} = \dfrac{256}{g}\)

ganeshie8 (ganeshie8):

cross multiply and solve \(\large g\)

OpenStudy (jocelynstalnaker):

1280/2 640=g <<< Is that right or no?

ganeshie8 (ganeshie8):

looks okay ^

ganeshie8 (ganeshie8):

\(\large \dfrac{2}{5} = \dfrac{256}{g}\) \(\large 2g = 1280\) \(\large g = 1280/2 = 640\)

OpenStudy (jocelynstalnaker):

Yes, So that would be my answer to E?

ganeshie8 (ganeshie8):

Yep !

OpenStudy (jocelynstalnaker):

Yay! Thank you so much!!

ganeshie8 (ganeshie8):

np :) whats ur opinion on last part ?

OpenStudy (jocelynstalnaker):

Oh wait! Didnt you say part B was incorrect?

ganeshie8 (ganeshie8):

yeah lets look at part B may be...

ganeshie8 (ganeshie8):

`Domain` is just a fancy word for `all the x values` that u can feed a function with

OpenStudy (jocelynstalnaker):

Yes, and i chose the number 5

ganeshie8 (ganeshie8):

5 is part of the domain, but its not `THE domain`

ganeshie8 (ganeshie8):

Your function from part A : g(x) = 128x

OpenStudy (jocelynstalnaker):

Yes

ganeshie8 (ganeshie8):

can you tell me what `x` represents here ?

OpenStudy (jocelynstalnaker):

For each cup?

ganeshie8 (ganeshie8):

yes x represents `cup of flour` can x be negative ? can x be 0 ? can x be positive ?

ganeshie8 (ganeshie8):

Notice that it makes no sense to say : use `-2` cups of flour for a recipe

OpenStudy (jocelynstalnaker):

No it cant be negative, well if it was 0 then why would the question say "flour" but it can be positive

ganeshie8 (ganeshie8):

however 0 is fine and it can be any positive value. So the domain is x >= 0

zepdrix (zepdrix):

They asked for a `reasonable domain`. Kendra can't bake infinite apple pies :D I think 5 cups is fine. I would have said 20 maybe, but whatev.

OpenStudy (jocelynstalnaker):

Oh ok

OpenStudy (jocelynstalnaker):

Greater than 0

ganeshie8 (ganeshie8):

0 is good to have - it represents the case when you feel lazy, change ur mind and give up on ur friends and decide not to make the pie

ganeshie8 (ganeshie8):

Hope you can answer the last part now, give it a try :)

OpenStudy (jocelynstalnaker):

Lol, ok so x>=0 is the answer and not just 5

OpenStudy (jocelynstalnaker):

I have F F- x represents the number of cups and a negative number -2 is not a reasonable in this situation because you cant have -2 cups.

zepdrix (zepdrix):

\(\Large\rm 0\lt x\lt 5\) is a reasonable domain. You should stick with the 5 in my opinion. :\

zepdrix (zepdrix):

Woops `or equal to` on the lower bound -_- my bad.

ganeshie8 (ganeshie8):

You have it !

OpenStudy (jocelynstalnaker):

Ok, Thank you guys so much for your help!

ganeshie8 (ganeshie8):

np, you're welcome :)